OpenStudy (anonymous):

142 2/9 sq yd = __ sq ft?

OpenStudy (jamesj):

Same principle as the last three problems you have posted. How many sq ft in 1 sq yard? (Other posters: please don't answer the answer yet ;-) )

OpenStudy (jamesj):

Let me ask you: how many feet in 1 yard?

OpenStudy (anonymous):

3 ft

OpenStudy (jamesj):

Right, so now, draw a diagram of 1 sq yard, and tell me how many sq feet there are in 1 sq yard

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

im terrible @ math

OpenStudy (jamesj):

Draw out 1 sq yard, and mark along the horizontal sides and the vertical sides the 3 feet Then draw a grid of the all the square feet in that one square yard

OpenStudy (anonymous):

its like u r speaking another language :<

OpenStudy (anonymous):

anyone answer plz.....

OpenStudy (jamesj):

give me a second, and I'll show you the diagram ... take a couple of minutes to draw

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

chef-dimples, try to understand what JamesJ is saying, really try.

OpenStudy (jamesj):

Ok. So this diagram shows one sq yard. Each side 1 yard long.

OpenStudy (jamesj):

And each yard comprises of 3 feet. So we have a grid of sq feet inside the sq yard. How many sq feet are there in this sq yard? Count them up

OpenStudy (anonymous):

3*4=12

OpenStudy (jamesj):

Seriously? Look again.

OpenStudy (anonymous):

or is it 9 since thats how many little boxes i see?

OpenStudy (jamesj):

Yes, 9. Or 3 times 3.

OpenStudy (jamesj):

So 1 sq yard = 9 sq feet

OpenStudy (anonymous):

ok so i divide 142 2/9 by 9?

OpenStudy (jamesj):

Think about it. Are you going to have more sq feet than sq yards or less sq feet than sq yards?

OpenStudy (jamesj):

If you divide you are going to have less. Can that possibly be right?

OpenStudy (anonymous):

ugh! maybe im not understanding this word problem i am reading!!!!

OpenStudy (jamesj):

If 1 sq yard = 9 sq feet then 10 sq yard = 90 sq feet and 100 sq yard = 900 sq feet So 142 2/9 sq yd = __ sq ft?

OpenStudy (anonymous):

its asking me to make the necessary conversion on how much carpet needs to be ordered if its sold by sq ft only!

OpenStudy (jamesj):

Yes, so what are we doing here. If we have a certain number of sq yards, how do we get the number in sq feet. Look at the examples above.

OpenStudy (anonymous):

multiply

OpenStudy (jamesj):

1 sq yard = 9 sq feet 5 sq yard = 45 sq feet 10 sq yard = 90 sq feet 100 sq yard = 900 sq feet

OpenStudy (jamesj):

Yes. Multiply.

OpenStudy (anonymous):

multiply the 142 by 9 then the 2/9 by 9?

OpenStudy (jamesj):

Sure

OpenStudy (anonymous):

142*9=1278 and 2/9*9=1.98 right?

OpenStudy (jamesj):

2/9 x 9 = 2 !!!!

OpenStudy (jamesj):

For example: \[\frac{2}{9} \times 9 = \frac{18}{2} = 9\]

OpenStudy (anonymous):

i never know when i should round off, omg, i suck @ math

OpenStudy (jamesj):

It's not a question of rounding. It's a basic question of manipulating fractions. If I have nine 9ths, i.e., 9 x (1/9), then that is 1. Take a pie, cut it up into 9 equal slices, how much are the nine slices? They are the whole pie.

OpenStudy (jamesj):

Anyway, what is the answer to 142 2/9 sq yd = __ sq ft?

OpenStudy (anonymous):

1287 sq ft

OpenStudy (anonymous):

1278 + 9 =1287?

OpenStudy (jamesj):

No. 142 x 9 = 1,278 2/9 x 9 = 2 So (142 2/9) x 9 = 142 x 9 + 2/9 x 9 = 1,278 + 2 = 1,280

OpenStudy (anonymous):

okay lord thank u! how am i going to make it through my mid terms in 2 weeks

OpenStudy (jamesj):

Practice, practice, practice. All of us have had to do it. You should go back to the questions you have posted earlier today and make sure you understand the answers.

OpenStudy (anonymous):

u r a life saver thank u